OpenStudy (anonymous):

starting with the graph f(x)=2^x; shifting f(x) 3 units to the right y would equal what?

OpenStudy (anonymous):

as f(x) = 2^x shifting it to the right 3 units you will get f(x-3) = 2^(x-3)

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

welcome